# CompBench: A Comparative Reasoning Benchmark for Multimodal LLMs

### v1

@@ Abstract
The ability to compare objects, scenes, or situations is crucial for effective decision-making and problem-solving in everyday life. For instance, comparing the freshness of apples enables better choices during grocery shopping, while comparing sofa designs helps optimize the aesthetics of our living space. Despite its significance, the comparative capability is largely unexplored in artificial general intelligence (AGI). In this paper, we introduce CompBench, a benchmark designed to evaluate the comparative reasoning capability of multimodal large language models (MLLMs). CompBench mines and pairs images through visually oriented questions covering eight dimensions of relative comparison: visual attribute, existence, state, emotion, temporality, spatiality, quantity, and quality. We curate a collection of around 40404040K image pairs using metadata from diverse vision datasets and CLIP similarity scores. These image pairs span a broad array of visual domains, including animals, fashion, sports, and both outdoor and indoor scenes. The questions are carefully crafted to discern relative characteristics between two images and are labeled by human annotators for accuracy and relevance. We use CompBench to evaluate recent MLLMs, including GPT-4V(ision), Gemini-Pro, and LLaVA-1.6. Our results reveal notable shortcomings in their comparative abilities. We believe CompBench not only sheds light on these limitations but also establishes a solid foundation for future enhancements in the comparative capability of MLLMs.
Figure 1: CompBench offers diverse triplets comprising two images, a question about their relativity, and an answer to cover eight types of relativity (see §1). See examples along with predictions of GPT-4V [1].
@@ 1 Introduction
In this paper, we introduce CompBench, a V&L benchmark dedicated to evaluating the comparative reasoning capabilities of MLLMs (Figure 1). CompBench comprises 39.839.839.839.8K triplets, each containing 1) a pair of visually or semantically relevant images 2) a question about their relativity, and 3) a ground-truth answer. We consider a wide range of questions categorized into eight aspects of relativity. Attribute Relativity tests the ability to recognize relative attributes [42] such as size, color, texture, shape, and pattern. For instance, given two images of birds, we ask MLLMs to compare the length of their beaks (e.g., “Which bird has longer beaks?”). Existential Relativity assesses the comprehension of existence in comparisons, asking questions like “Which trait is in the left butterfly but not in the right butterfly?”. State/Emotion Relativity examines if MLLMs can identify state variations, such as different degrees of baking and smiling. Temporal Relativity evaluates the understanding of time-related changes between two objects or scenes (e.g., “Which video frame happens earlier during a free kick?”). Spatial Relativity checks the ability to tell spatial differences (e.g., “Which cup looks further?”). Finally, Quantitiy/Quality Relativity investigates whether an MLLM understands the relativity of quantity and quality (e.g., “Which image contains more animal instances?”).
We systematically benchmark representative MLLMs on CompBench, including GPT-4V [1], Gemini1.0-Pro [54], LLaVA-1.6 [32], and VILA-1.5 [31]. Specifically, we concatenate two images horizontally (i.e., left and right) as the visual input. We then prompt MLLMs to answer questions about the relativity between these two images. When applicable, we also investigate a two-stage reasoning strategy, starting by asking a refined question about each image independently (e.g., “How many animal instances are in the image?”), followed by a pure language question (e.g., “Based on the descriptions, which image has more animal instances?”). Our results reveal notable shortcomings in existing MLLMs’ comparative abilities, especially in Existence, Spatiality, and Quantity Relativity. We conduct further analyses of error cases, offering insights for future MLLMs’ improvements.
In sum, CompBench has several advantages: (i) CompBench introduces new perspectives to evaluate MLLMs — comparative reasoning capabilities about relativity. (ii) CompBench provides extensive coverage across eight relativities and fourteen domains. (iii) CompBench benchmarks recent MLLMs, accompanied by detailed analyses and insights for future improvement. (iv) CompBench is extensible — we identify multiple data sources that can be further incorporated.
@@ 2 Related Work
MLLM benchmarks. Earlier, MLLMs were evaluated on traditional V&L tasks, such as visual question answering (VQA) [21, 22, 49], image captioning [10, 2], and image-text retreival [30, 11]. Recently, a range of new and intriguing V&L tasks [36, 52] have emerged to assess MLLMs’ capabilities across various dimensions. These include comprehension and reasoning about charts [37], diagrams [38], scene text [50, 39], web navigation [16], expert-level multimodal understanding [60], etc. Our CompBench complements these efforts by focusing on a new dimension, MLLMs’ comparative reasoning capacity on a pair of visually or semantically relevant images.
Multi-image datasets. Several existing datasets [42, 53, 17, 25, 62, 27] provide multi-image data (e.g., pairs of images), but they serve different purposes (e.g., not for evaluating MLLMs) or have relatively limited scopes. NLVR2 [53] labels each image pair with a caption that may or may not be relevant to the images, asking models to predict the caption’s relevance (i.e., image-text matching). A few datasets [27, 7, 61] synthesize multi-image data for instruction tuning (e.g., image editing). More relevant to ours are [17, 25, 62, 42]. Birds-to-Words [17] aims to describe the difference between two birds; Sopt-the-diff [25] focuses on the difference between two outdoor scenes; Q-bench2 [62] compares the quality (e.g., blurriness) between two images; Relative Attributes [42] compares the relativeness of attributes between two facial or natural images. However, these datasets have limited scopes, only targeting specific domains or questions. In contrast, our CompBench defines eight relative comparisons, covering a wide range of relativities in the real world. Our image pairs are curated from fourteen diverse visual domains. We believe this offers the V&L community a more comprehensive benchmark to assess the comparative capabilities of current leading MLLMs.
Learning to rank & learning with preference. Several research topics are relevant to ours and may benefit from our CompBench. Learning to rank (LTR) [29, 33, 8] aims to realize a scoring function that can rank examples (e.g., images) based on certain aspects, such as facial ages [40, 9] and degrees of attributes’ presence [42]. Typically, an LTR model takes one example as input; the model is trained with pairs of examples such that the output scores match the ground-truth orders. Recently, learning with preference information [18] has become a mainstream approach to fine-tuning LLMs for alignment [47, 12]. Unlike our focus, these works usually collect pairs of outputs (e.g., answers to a question) with humans’ preferences to supervise model fine-tuning.
@@ 3 Why Do We Study Comparative Reasoning?
4 CompBench Benchmark
We introduce CompBench, a multimodal benchmark designed to assess the comparative reasoning abilities of MLLMs across various dimensions. In what follows, we first describe the types of comparative capabilities that CompBench aims to evaluate (§4.1). Next, we outline our methodology for collecting images, followed by how we annotate associated questions and answers to evaluate these capabilities (§4.2). Lastly, we provide detailed statistics on CompBench and discuss its data quality (§4.3). Figure 2 illustrates the overall pipeline used to develop CompBench.
Figure 2: CompBench curation pipeline, including data selection, question generation, answer annotation, and verification. We rely on combinations of humans, computer programs, MLLMs (specifically GPT-4V [1]), and CLIP similarity [46] to select images and generate questions, based on relativity types and available metadata.
@@ 4.2 Dataset Curation
One major challenge in constructing CompBench is mining image pairs that reflect the aforementioned relativities. Fortunately, many publicly accessible datasets in vision and V&L offer detailed annotations and metadata. We carefully investigate these datasets and identify a seed set of fourteen datasets that align with the eight relativity types (§4.1), covering a wide range of domains like open-domain, fashion, animal, sports, automotive, facial, and both outdoor and indoor scenes (cf. Right in Table 1). Below, we outline the datasets for each relativity type and the process for generating triplets of image pairs, a question, and an answer. Please see the supplementary material for details.
@@ 4.2.6 Spatiality
Table 1: Overall statistics of CompBench.
@@ 4.3 Quality Control and Dataset Statistics
To ensure the integrity of CompBench, we ask annotators to exclude poor-quality examples, such as those with low-resolution images or questions that are irrelevant or nonsensical about the images. The annotators also filter out image pairs with ambiguous relativities, for example, image pairs with indistinguishable smiling degrees. To faithfully assess fine-grained capabilities, we also apply the CLIP visual similarity to Existence, removing image pairs with salient differences. Additionally, we implement a rigorous cross-verification process, where each annotator confirms the accuracy of others’ answers. Only samples that receive unanimous approval from annotators are kept. Consequently, our CompBench benchmark comprises 39.8K diverse triplets (eight relativities from fourteen visual domains) with high quality and reliability. Please see Table 1 for the statistics.
Human Annotators & Evaluators. We recruited five in-house human annotators from our research team to work on CompBench. The annotators are instructed to avoid generating any personally identifiable information or offensive content during the annotation process. Furthermore, we recruited another five human evaluators, who were not involved in the annotation, to measure the upper bound model performance on CompBench (Table 4). The workloads for annotation and evaluation were distributed equally among annotators and evaluators.
@@ 5.1 Experimental Setup
Baselines. We use our CompBench to evaluate several leading MLLMs. This includes two powerful proprietary models, GPT-4V(ision) [1] and Gemini1.0-Pro [^1] Due to limited public testing quota available for Gemini-1.5 during our study, we opted for Gemini-1.0 Pro. [54], and two open-source alternatives, LLaVA-1.6 [32] and VILA-1.5 [31]. GPT-4V(ision) and Gemini excel in various vision and language tasks, such as VQA [21], OCR interpretation [39], spatial reasoning [37], and college-level subject knowledge [60]. LLaVA-1.6 and VILA-1.5 also demonstrate competitive performance against these proprietary giants on some tasks. Our focus is to investigate whether these cutting-edge models can extend their capabilities to the realm of multi-image relative comparison. We evaluate proprietary models via their official APIs and open-source models using (or fine-tuning on) NVIDIA RTX 6000 Ada GPUs. For more details, please refer to the supplementary material.
Evaluation tasks & metrics. We divide our CompBench into a test split (31.8K) and a held-out split (7.9K), using an 80:20 ratio. The latter is reserved for future developments (e.g., prompt engineering). By default, we concatenate the image pairs horizontally (i.e., left and right) as the visual input to MLLMs, and prompt MLLMs to answer questions about the relativity between these images. To facilitate automated evaluation, we include the possible answers as options in the questions. For Existence and Quality, there are multiple options (typically more than two). For Quantity, there are three options: left/right/same. For other types, there are binary options: left/right. We employ the standard accuracy as our evaluation metric. A question is answered correctly if the model prediction exactly matches the ground-truth answer. Further details are included in the supplementary material.
@@ 5.3 Further Analyses
Human evaluation. We investigate how much current MLLMs (e.g., GPT-4V [1]) lag behind human performance. We conduct a preliminary human evaluation using $140140$ examples randomly sampled from the sixteen tasks (columns) in Table 1. We ask five human evaluators, different from our annotators, to answer these questions and average their performance. As shown in Table 4, the performance of GPT-4V on these examples is approximately 18% below that of humans. This not only highlights the challenge of our CompBench but also underscores the limited capabilities of current MLLMs in multi-image relative comparison.
6 Conclusion
We introduce CompBench, a comprehensive benchmark designed to evaluate comparative reasoning in multimodal LLMs (MLLMs). CompBench offers extensive coverage of eight relative comparisons between pairs of images drawn from fourteen diverse domains. CompBench evaluates recent MLLMs, offering detailed analyses and insights for future advancements.
@@ Acknowledgments
This research is supported in part by grants from the National Science Foundation (IIS-2107077, OAC-2112606, OAC-2118240). We would like to extend our thanks to our colleagues from the OSU MLB lab for their thoughtful feedback and discussions. In addition, we are grateful for the generous support of computational resources provided by the Ohio Supercomputer Center.
@@
All codes, data, and instructions for our CompBench can be found in https://github.com/RaptorMai/CompBench. CompBench is released under a Creative Commons Attribution 4.0 License (CC BY 4.0).
@@
  • Appendix B: CompBench curation and model evaluation (cf. §4.2 and §5.1 in the main text).

### v2

# MLLM-CompBench: A Comparative Reasoning Benchmark for Multimodal LLMs

@@ Abstract
The ability to compare objects, scenes, or situations is crucial for effective decision-making and problem-solving in everyday life. For instance, comparing the freshness of apples enables better choices during grocery shopping, while comparing sofa designs helps optimize the aesthetics of our living space. Despite its significance, the comparative capability is largely unexplored in artificial general intelligence (AGI). In this paper, we introduce MLLM-CompBench, a benchmark designed to evaluate the comparative reasoning capability of multimodal large language models (MLLMs). MLLM-CompBench mines and pairs images through visually oriented questions covering eight dimensions of relative comparison: visual attribute, existence, state, emotion, temporality, spatiality, quantity, and quality. We curate a collection of around 40404040K image pairs using metadata from diverse vision datasets and CLIP similarity scores. These image pairs span a broad array of visual domains, including animals, fashion, sports, and both outdoor and indoor scenes. The questions are carefully crafted to discern relative characteristics between two images and are labeled by human annotators for accuracy and relevance. We use MLLM-CompBench to evaluate recent MLLMs, including GPT-4V(ision), Gemini-Pro, and LLaVA-1.6. Our results reveal notable shortcomings in their comparative abilities. We believe MLLM-CompBench not only sheds light on these limitations but also establishes a solid foundation for future enhancements in the comparative capability of MLLMs.
Figure 1: MLLM-CompBench offers diverse triplets comprising two images, a question about their relativity, and an answer to cover eight types of relativity (see §1). See examples along with predictions of GPT-4V [1].
@@ 1 Introduction
In this paper, we introduce MLLM-CompBench, a V&L benchmark dedicated to evaluating the comparative reasoning capabilities of MLLMs (Figure 1). MLLM-CompBench comprises 39.839.839.839.8K triplets, each containing 1) a pair of visually or semantically relevant images 2) a question about their relativity, and 3) a ground-truth answer. We consider a wide range of questions categorized into eight aspects of relativity. Attribute Relativity tests the ability to recognize relative attributes [44] such as size, color, texture, shape, and pattern. For instance, given two images of birds, we ask MLLMs to compare the length of their beaks (e.g., “Which bird has longer beaks?”). Existential Relativity assesses the comprehension of existence in comparisons, asking questions like “Which trait is in the left butterfly but not in the right butterfly?”. State/Emotion Relativity examines if MLLMs can identify state variations, such as different degrees of baking and smiling. Temporal Relativity evaluates the understanding of time-related changes between two objects or scenes (e.g., “Which video frame happens earlier during a free kick?”). Spatial Relativity checks the ability to tell spatial differences (e.g., “Which cup looks further?”). Finally, Quantitiy/Quality Relativity investigates whether an MLLM understands the relativity of quantity and quality (e.g., “Which image contains more animal instances?”).
We systematically benchmark representative MLLMs on MLLM-CompBench, including GPT-4V [1], Gemini1.0-Pro [58], LLaVA-1.6 [33], and VILA-1.5 [32]. Specifically, we concatenate two images horizontally (i.e., left and right) as the visual input. We then prompt MLLMs to answer questions about the relativity between these two images. When applicable, we also investigate a two-stage reasoning strategy, starting by asking a refined question about each image independently (e.g., “How many animal instances are in the image?”), followed by a pure language question (e.g., “Based on the descriptions, which image has more animal instances?”). Our results reveal notable shortcomings in existing MLLMs’ comparative abilities, especially in Existence, Spatiality, and Quantity Relativity. We conduct further analyses of error cases, offering insights for future MLLMs’ improvements.
In sum, MLLM-CompBench has several advantages: (i) MLLM-CompBench introduces new perspectives to evaluate MLLMs — comparative reasoning capabilities about relativity. (ii) MLLM-CompBench provides extensive coverage across eight relativities and fourteen domains. (iii) MLLM-CompBench benchmarks recent MLLMs, accompanied by detailed analyses and insights for future improvement. (iv) MLLM-CompBench is extensible — we identify multiple data sources that can be further incorporated. Remark. During the conference, we found a concurrent work by Kazemi et al. [28] that also studies MLLM’s reasoning capability with multiple images, specifically focusing on math, physics, logic, code, table/chart understanding, spatial and temporal domains. We encourage readers to consult their work as well.
@@ 2 Related Work
MLLM benchmarks. Earlier, MLLMs were evaluated on traditional V&L tasks, such as visual question answering (VQA) [21, 22, 53], image captioning [10, 2], and image-text retreival [31, 11]. Recently, a range of new and intriguing V&L tasks [37, 56] have emerged to assess MLLMs’ capabilities across various dimensions. These include comprehension and reasoning about charts [38], diagrams [39], scene text [54, 40], web navigation [16], expert-level multimodal understanding [65], etc. Our MLLM-CompBench complements these efforts by focusing on a new dimension, MLLMs’ comparative reasoning capacity on a pair of visually or semantically relevant images.
Multi-image datasets. Several existing datasets [44, 57, 17, 25, 67, 27] provide multi-image data (e.g., pairs of images), but they serve different purposes (e.g., not for evaluating MLLMs) or have relatively limited scopes. NLVR2 [57] labels each image pair with a caption that may or may not be relevant to the images, asking models to predict the caption’s relevance (i.e., image-text matching). A few datasets [27, 7, 66] synthesize multi-image data for instruction tuning (e.g., image editing). More relevant to ours are [17, 25, 67, 44]. Birds-to-Words [17] aims to describe the difference between two birds; Sopt-the-diff [25] focuses on the difference between two outdoor scenes; Q-bench2 [67] compares the quality (e.g., blurriness) between two images; Relative Attributes [44] compares the relativeness of attributes between two facial or natural images. However, these datasets have limited scopes, only targeting specific domains or questions. In contrast, our MLLM-CompBench defines eight relative comparisons, covering a wide range of relativities in the real world. Our image pairs are curated from fourteen diverse visual domains. We believe this offers the V&L community a more comprehensive benchmark to assess the comparative capabilities of current leading MLLMs.
Learning to rank & learning with preference. Several research topics are relevant to ours and may benefit from our MLLM-CompBench. Learning to rank (LTR) [30, 34, 8] aims to realize a scoring function that can rank examples (e.g., images) based on certain aspects, such as facial ages [41, 9] and degrees of attributes’ presence [44]. Typically, an LTR model takes one example as input; the model is trained with pairs of examples such that the output scores match the ground-truth orders. Recently, learning with preference information [18] has become a mainstream approach to fine-tuning LLMs for alignment [50, 12]. Unlike our focus, these works usually collect pairs of outputs (e.g., answers to a question) with humans’ preferences to supervise model fine-tuning.
@@ 3 Why Do We Study Comparative Reasoning?
4 MLLM-CompBench Benchmark
We introduce MLLM-CompBench, a multimodal benchmark designed to assess the comparative reasoning abilities of MLLMs across various dimensions. In what follows, we first describe the types of comparative capabilities that MLLM-CompBench aims to evaluate (§4.1). Next, we outline our methodology for collecting images, followed by how we annotate associated questions and answers to evaluate these capabilities (§4.2). Lastly, we provide detailed statistics on MLLM-CompBench and discuss its data quality (§4.3). Figure 2 illustrates the overall pipeline used to develop MLLM-CompBench.
Figure 2: MLLM-CompBench curation pipeline, including data selection, question generation, answer annotation, and verification. We rely on combinations of humans, computer programs, MLLMs (specifically GPT-4V [1]), and CLIP similarity [49] to select images and generate questions, based on relativity types and available metadata.
@@ 4.2 Dataset Curation
One major challenge in constructing MLLM-CompBench is mining image pairs that reflect the aforementioned relativities. Fortunately, many publicly accessible datasets in vision and V&L offer detailed annotations and metadata. We carefully investigate these datasets and identify a seed set of fourteen datasets that align with the eight relativity types (§4.1), covering a wide range of domains like open-domain, fashion, animal, sports, automotive, facial, and both outdoor and indoor scenes (cf. Right in Table 1). Below, we outline the datasets for each relativity type and the process for generating triplets of image pairs, a question, and an answer. Please see the supplementary material for details.
@@ 4.2.6 Spatiality
Table 1: Overall statistics of MLLM-CompBench.
@@ 4.3 Quality Control and Dataset Statistics
To ensure the integrity of MLLM-CompBench, we ask annotators to exclude poor-quality examples, such as those with low-resolution images or questions that are irrelevant or nonsensical about the images. The annotators also filter out image pairs with ambiguous relativities, for example, image pairs with indistinguishable smiling degrees. To faithfully assess fine-grained capabilities, we also apply the CLIP visual similarity to Existence, removing image pairs with salient differences. Additionally, we implement a rigorous cross-verification process, where each annotator confirms the accuracy of others’ answers. Only samples that receive unanimous approval from annotators are kept. Consequently, our MLLM-CompBench benchmark comprises 39.8K diverse triplets (eight relativities from fourteen visual domains) with high quality and reliability. Please see Table 1 for the statistics.
Human Annotators & Evaluators. We recruited five in-house human annotators from our research team to work on MLLM-CompBench. The annotators are instructed to avoid generating any personally identifiable information or offensive content during the annotation process. Furthermore, we recruited another five human evaluators, who were not involved in the annotation, to measure the upper bound model performance on MLLM-CompBench (Table 4). The workloads for annotation and evaluation were distributed equally among annotators and evaluators.
@@ 5.1 Experimental Setup
Baselines. We use our CompBench  [^1] We use CompBench and MLLM-CompBench interchangeably. to evaluate several leading MLLMs. This includes two powerful proprietary models, GPT-4V(ision) [1] and Gemini1.0-Pro [^2] Due to limited public testing quota available for Gemini1.5 during our study, we opted for Gemini1.0-Pro. [58], and two open-source alternatives, LLaVA-1.6 [33] and VILA-1.5 [32]. GPT-4V(ision) and Gemini excel in various vision and language tasks, such as VQA [21], OCR interpretation [40], spatial reasoning [38], and college-level subject knowledge [65]. LLaVA-1.6 and VILA-1.5 also demonstrate competitive performance against these proprietary giants on some tasks. Our focus is to investigate whether these cutting-edge models can extend their capabilities to the realm of multi-image relative comparison. We evaluate proprietary models via their official APIs and open-source models using (or fine-tuning on) NVIDIA RTX 6000 Ada GPUs. For more details, please refer to the Appendix C in supplementary material.
Evaluation tasks & metrics. We divide our CompBench into a test split (31.8K) and a held-out split (7.9K), using an 80:20 ratio. The latter is reserved for future developments (e.g., prompt engineering). By default, we concatenate the image pairs horizontally (i.e., left and right) as the visual input to MLLMs, and prompt MLLMs to answer questions about the relativity between these images. To facilitate automated evaluation, we include the possible answers as options in the questions. For Existence and Quality, there are multiple options (typically more than two). For Quantity, there are three options: left/right/same. For other types, there are binary options: left/right. We employ the standard accuracy as our evaluation metric. A question is answered correctly if the model prediction exactly matches the ground-truth answer. Further details are included in the Appendix B.
@@ 5.3 Further Analyses
Human evaluation. We investigate how much current MLLMs (e.g., GPT-4V) lag behind human performance. We conduct a preliminary human evaluation using $140140$ examples randomly sampled from the sixteen tasks (columns) in Table 1. We ask five human evaluators, different from our annotators, to answer these questions and average their performance. As shown in Table 4, the performance of GPT-4V on these examples is approximately 18% below that of humans. This not only highlights the challenge of our CompBench but also underscores the limited capabilities of current MLLMs in multi-image relative comparison.
5.4 Evaluation of Recent MLLMs Released After the NeurIPS Deadline
Since our paper submission in early June 2024, several new MLLMs have been released, such as GPT-4o [42] or Gemini1.5-Pro [52]. In Table 5, we present a comparative analysis of GPT-4V with the recently released GPT-4o, alongside Gemini1.0-Pro with Gemini1.5-Pro. These upgraded models demonstrate substantial improvements over their previous versions, with GPT-4o showing marked gains in existence (MB and SD) and spatial (ND) relativities, while Gemini1.5-Pro achieves broader enhancements across multiple relational dimensions. This progress likely results from enhanced training approaches, including scaling model and data along with refined learning strategies. Investigating exactly how these methods drive GPT-4o’s substantial gains on our benchmark could be a valuable direction for future research. Nonetheless, we note that the performance of GPT-4o remains mediocre in several relativities, such as spatiality and quantity.
Model Attribute Exist. State Emot. Temp. Spat. Quan. Qual. Avg
ST FA VA CU WF MB SD ST VA CE FE SN $C$ ND VQ QB
GPT-4o 92.3 97.0 86.3 74.7 84.5 81.2 67.2 95.8 89.6 96.6 91.1 72.0 83.3 68.2 67.8 81.2 83.1
Improvement 0.5 8.0 9.4 3.3 12.4 22.9 25.3 3.6 1.8 4.8 7.7 0.6 9.6 12.1 4.0 8.2 8.4
Gemini1.5-Pro 79.2 91.8 77.7 71.4 72.8 55.4 58.7 91.0 84.0 93.0 87.3 50.3 70.3 68.3 64.8 70.5 74.2
Improvement 7.3 15.5 8.4 11.5 17.9 1.7 5.7 9.2 13.3 32.4 16.1 -4.8 12.1 11.7 10.2 11.0 11.2
Table 5: Results of new MLLM models (GPT4-o and Gemini1.5-Pro) released after NeurIPS deadline. The top-performing model in each task is indicated in bold. Both upgraded MLLM models (GPT-4o & Gemini1.5-Pro) exhibit significant improvements over their previous versions (GPT-4V & Gemini1.0-Pro).
6 Conclusion and Future Work
In this work, we introduce MLLM-CompBench, a comprehensive benchmark designed to evaluate comparative reasoning in multimodal LLMs (MLLMs), offering detailed analyses and insights for future advancements. As future work, we plan to incorporate more challenging datasets into each type of relative comparison in MLLM-CompBench. For instance, additional video datasets could be explored for temporal relativity, such as cooking activities [68] or other sports [62, 48]. Moreover, expanding the scope of comparative reasoning relativities holds promise. Examples include similarity comparisons (e.g., “Identify similar objects between the two images.”) and comparisons involving more than two images (e.g., “Given images showing various views of one object along with a few of a different object, the model should identify the outliers.”). We envision that MLLM-CompBench will serve as a valuable tool, paving the way for advancing comparative reasoning in MLLMs.
@@ Acknowledgments
This research is supported in part by grants from the National Science Foundation (IIS-2107077, OAC-2112606, OAC-2118240). We extend our sincere appreciation to our colleagues from the OSU MLB lab—Jinsu Yoo, Ping Zhang, Jiaman Wu, Ziwei Li, and Tai-Yu Pan—for their thoughtful feedback and discussions. Finally, we are grateful for the generous support of computational resources provided by the Ohio Supercomputer Center.
@@
All codes, data, and instructions for our MLLM-CompBench can be found in https://github.com/RaptorMai/CompBench. MLLM-CompBench is released under a Creative Commons Attribution 4.0 License (CC BY 4.0).
@@
  • Appendix B: MLLM-CompBench curation and model evaluation (cf. §4.2 and §5.1 in the main text).